OpenStudy (anonymous):

Help plz Represents this function: \[y=2x^{3}-3x ^{2}-12x+8\]

OpenStudy (anonymous):

@sami-21

OpenStudy (helder_edwin):

do u want the graph of this function? is this what u mean by "represent"??

OpenStudy (anonymous):

i am thinking the same. what does this "represents" mean here.

OpenStudy (anonymous):

graph

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first this is cubic polynomial its general graph is  put x=0 to find where it crosses y axis when x=0 y=8 it means graph crosses y axis at y=8 so it should be like this

OpenStudy (anonymous):

Graph draw karnay kay liye kuch x points select karo like x=0,1,2,3,4 inn kay liye y value find karo by putting these values in the function for example x=0 y=8 we have one point now (0,8) x=1 y=-5 (1,-5) x=2 y=-12 (2,-12) x=3 y=-1 (3,-1) x=4 y=40 (4,40) now we have four points (1,-5),(1,-5),(2,-12),(3,-1),(4,40) inn ko draw karo paper par kar lo gi ?

OpenStudy (anonymous):

ok .. let me try

OpenStudy (anonymous):

@sami-21 for that eq can we use that approach: 'The slope of the tangent line is the value of the derivative. Horizontal lines have a slope of m=0 then find 'x' while m=derivative of function w.r.t x"

OpenStudy (anonymous):

 ye may nay draw kiye hain roughly.